### v1

# The role of charm and unflavored mesons in prompt atmospheric lepton fluxes

Abstract
The all-sky very-high-energy ($10^{4}-10^{6}$ GeV) atmospheric muon flux measured by IceCube shows a spectral hardening at the highest energies, indicating the presence of a prompt component. IceCube has also measured the atmospheric muon neutrino flux at high energy. However, since this flux is dominated by astrophysical neutrinos, only an upper bound can be placed on the prompt atmospheric $\nu_{\mu}+\bar{\nu}_{\mu}$ contribution. In this work, we provide a new evaluation of the prompt atmospheric muon flux including an intrinsic charm component in the cosmic ray-air interactions. The latter enhances the forward production of $\bar{D}^{0}$, $D^{-}$, and $\Lambda_{c}$, which subsequently decay into final states containing muons and muon neutrinos. We show how the increase in the prompt muon flux due to intrinsic charm is accompanied by a corresponding enhancement in the prompt muon neutrino flux. We implement different intrinsic charm production models in MCEq to calculate the resulting lepton fluxes. We discuss the challenges of achieving predictions that are simultaneously consistent with both IceCube’s high-energy atmospheric muon flux measurements and IceCube upper bound on the prompt muon neutrino flux, and we quantify the resulting discrepancies. As possible solutions, we explore scaling of the unflavored meson contributions to the prompt atmospheric muon flux to assess how such adjustments can reconcile these differences. The tensions emphasized in our work call for a refinement of the hadronic interaction models, especially the production of unflavored mesons, and for new experimental data sensitive to unflavored meson and heavy flavor production with reliable estimates of the associated uncertainties. We suggest that the energy and zenith angle dependence of muon and neutrino flux ratios from future neutrino telescope measurements may help to disentangle different scenarios.
preprint: DESY-25-194
1 Introduction
When high-energy cosmic rays enter the Earth’s atmosphere, they interact with air nuclei and initiate a cascade of secondary particles, including protons, neutrons, pions, kaons, and charm hadrons. These secondary hadrons undergo further interactions and decays, giving rise to the atmospheric fluxes of muons and neutrinos (see, e.g., ref. Gaisser:2002jj and references therein). Since the primary cosmic ray spectrum spans many orders of magnitude in energy, the secondary particles are also produced in a wide energy range Gaisser:2016uoy ; Lipari:1993hd . Among these, high-energy atmospheric leptons ($E>10^{4}$ GeV), in particular muons and neutrinos, are of central importance in astroparticle physics. The latter form a major background for neutrino telescopes searching for astrophysical neutrinos IceCube:2013low ; IceCube:2016umi ; IceCube:2020acn ; IceCube:2024fxo ; IceCube:2025dlr , making precise knowledge of the atmospheric neutrino fluxes essential. Such measurements, together with the complementary input from atmospheric muon fluxes, not only help testing and constraining hadronic interaction models (see, e.g., Fedynitch:2018cbl ; Fedynitch:2022vty ; Yanez:2023lsy ), but also enable robust searches for astrophysical neutrino signals.
Atmospheric leptons are produced through a chain of interactions and decays: cosmic rays interact with air nuclei with average atomic mass number of $\langle A_{\rm air}\rangle=14.5$ to produce hadrons. Muons and neutrinos come from leptonic or semi-leptonic decays of these hadrons. To evaluate the atmospheric lepton flux, incident cosmic ray nuclei of mass number $A_{N}$, charge $Z$ and energy $E_{0}$ are treated in the superposition approximation, i.e., the nucleus is treated as an ensemble of $Z$ protons and $A_{N}-Z$ neutrons, each carrying energy $E_{0}/A_{N}$ Gaisser:2016uoy . The resulting all-nucleon cosmic ray spectrum at the top of the atmosphere is denoted by $\phi^{0}_{N}(E)$. Although the cosmic ray composition above $E_{0}\simeq 10^{5}-10^{6}$ GeV Albrecht:2021cxw has been measured only indirectly, there is reasonably good agreement among cosmic ray composition models that $\phi^{0}_{N}(E)\sim E^{-\gamma}$ with the value of $\gamma$ approximately constant but different in different energy ranges (e.g., broken power law) Gaisser:2011klf ; Gaisser:2013bla ; Hoerandel:2002yg .
The magnitudes of the atmospheric lepton fluxes from hadron decays depend on the competition between hadron-air interactions and hadron decays. If the hadron interaction length is much longer than its decay length, the lepton flux at energy $E_{\ell}$ approximately follows the primary cosmic ray nucleon spectrum, scaling as $\sim E_{\ell}^{-\gamma}$. Conversely, when hadrons are more likely to interact than decay, the corresponding atmospheric lepton flux is suppressed, scaling as $E_{\ell}^{-\gamma-1}$ Lipari:1993hd . For the energies considered here, $E_{\ell}\gtrsim 10^{4}$ GeV, the contributions from $\pi^{\pm}$ and $K$ decays to muon and neutrino fluxes scale like $\sim E_{\ell}^{-\gamma-1}$. The combination of these two contributions is referred to as the conventional atmospheric lepton flux. In contrast, hadrons with much shorter lifetimes, such as charm hadrons, decay promptly over a wide energy range. The prompt atmospheric lepton flux from charm hadron decays scales as $\sim E_{\ell}^{-\gamma}$, up to lepton energies $E_{\ell}\sim 2\times 10^{7}$ GeV Gondolo:1995fq . In addition to charm hadron decays, prompt muon production also receives contributions from electromagnetic decays of light unflavored mesons such as $\eta,\ \eta^{\prime},\ \rho^{0},\ \omega,$ and $\phi$, despite their small branching fractions ($\sim 10^{-5}$$10^{-4}$) Illana:2010gh . Their extremely short lifetimes ensure decay before interaction, enhancing the prompt atmospheric muon flux at high energies by a factor of $\sim 2$ relative to the prompt neutrino flux.
[…]
Figure 1: The atmospheric ($\mu^{+}+\mu^{-}$) flux and atmospheric ($\nu_{\mu}+\bar{\nu}_{\mu}$) flux, both scaled by $E^{3}$ and angle-averaged over zenith angles less than 60, from MCEq Fedynitch:2018cbl using H3a cosmic ray flux model Gaisser:2011klf and Sibyll-2.3c Riehn:2017mfm ; Fedynitch:2018cbl for interactions. The dashed lines show the prompt contributions, and the solid lines show the total contributions (conventional + prompt).
Prompt neutrinos from charm decays in the atmosphere are often considered an irreducible background to astrophysical neutrinos since both fluxes are nearly isotropic at Earth. Their spectral behaviors differ, offering possible means of discrimination between the prompt atmospheric background and astrophysical neutrinos. In this paper, we focus on using the atmospheric muon flux as a complementary probe to atmospheric neutrinos. While many previous studies have focused primarily on neutrino fluxes Laha:2016dri ; Halzen:2016thi ; Ostapchenko:2022thy ; Garzelli:2023jlq , our analysis expands to include both muons and neutrinos (see also, ref. Das:2025snq ). As we describe in the next section, measurements of the atmospheric muon flux, particularly in the energy range where the prompt muon flux begins to dominate, exceed theory predictions, although with large experimental error bars Soldin:2023lbr ; Soldin:2018vak ; IceCube:2015wro . To solve this discrepancy, we first consider potential contributions from intrinsic charm. The intrinsic charm contributions are also constrained by prompt neutrino measurements. Therefore, with existing upper bounds on the prompt atmospheric muon neutrino flux from IceCube IceCube:2016umi ; IceCube:2015wro , we also assess the extent to which the contribution of unflavored mesons to the prompt atmospheric muon flux may be enhanced.
The paper is organized as follows: in Section 2, we discuss the cascade equations that describe the development of the atmospheric lepton fluxes as cosmic rays and the hadrons they produce propagate through the atmosphere. We also show results from the cascade equations for muons and muon neutrinos, along with IceCube data and upper bounds on the prompt neutrino flux, respectively. In Section 3, we introduce the framework we use for incorporating intrinsic charm effects. Section 4 presents our results, followed by a discussion and our conclusions in Section 5. For completeness, the Appendix includes details of the charm particle decays to leptons, needed to evaluate the lepton fluxes.
2 Atmospheric lepton fluxes
2.1 Cascade equations
Quantitative evaluations of the atmospheric flux begin with the incident all-nucleon cosmic ray spectrum at the top of the atmosphere, $\phi_{N}^{0}(E)$. The flux per solid angle $\phi_{N}^{0}$ is isotropic at high energies. The cascade equations that describe the development of particle fluxes with subsequent interactions and decays Lipari:1993hd ; Gondolo:1995fq . The cascade equations are evaluated as a function of column depth $X$ which depends on the atmospheric density ($\rho$), path length ($l$) traversed by the particle in the atmosphere, altitude ($h$) and zenith angle ($\theta$),
$X(l,\theta)=\int_{l}^{\infty}dl^{\prime}\ \rho(h(l^{\prime},\theta))\,.$ (1)
The cascade equation for nucleons is
$\frac{d\phi_{N}(E,X)}{dX}$ $=$ $-\frac{\phi_{N}(E,X)}{\lambda_{N}(E)}+S(N\to N;E,X)\,$, (2)
where the source term for nucleon-nucleon regeneration is determined by
$S(k\to j;E,X)=\int_{E}^{\infty}dE^{\prime}\frac{\phi_{k}(E^{\prime},X)}{\lambda_{k}(E^{\prime})}\frac{dn(k\to j;E^{\prime},E)}{dE}\,$, (3)
for $k=j=N$. The energy distribution of outgoing particle $j$ with incoming particle $k$ is given by
$\frac{dn(k\to j;E^{\prime},E)}{dE}=\frac{1}{\sigma_{kA}(E^{\prime})}\frac{d\sigma(k\to j;E^{\prime},E)}{dE}\,.$ (4)
where $E^{\prime}$ and $E$ are energies of incoming particle $k$ and outgoing particle $j$, respectively. The interaction length depends on the cross section, with $\lambda_{k}=1/(N_{A}\sigma_{kA})$. For other hadrons, additional loss in flux occurs from decays, so
$\frac{d\phi_{j}(E,X)}{dX}=-\frac{\phi_{j}(E,X)}{\lambda_{j}(E)}-\frac{\phi_{j}(E,X)}{\lambda_{j}^{\rm dec}(E)}+\sum_{k}S(k\to j;E,X)\,$, (5)
where the hadron decay length is $\lambda^{\rm dec}_{j}(E)=\rho(h)\gamma_{j}(E)c\tau_{j}$ and $\gamma_{j}(E)=E/m_{j}$.
Muons and neutrinos are produced by hadron decays in the atmosphere (neglecting the photon conversion to $\mu^{+}+\mu^{-}$ processes, discussed below). Their cascade equations are
$\frac{d\phi_{\nu}(E,X)}{dX}$ $=$ $\sum_{j}S(j\to\nu;E,X)\,$ (6)
$\frac{d\phi_{\mu}(E,X)}{dX}$ $=$ $-\frac{\partial}{\partial E}\Bigl(\beta(E)\phi_{\mu}(E,X)\Bigr)+\sum_{j}S(j\to\mu;E,X)\,.$ (7)
The muon cascade equation has a term $\beta(E)\equiv dE/dX$ that accounts for electromagnetic energy loss due to ionization and radiative processes. The source terms for decays to muons and neutrinos involve the decay distributions rather than the differential cross sections, and the decay lengths rather than the interaction lengths.
The total column depth traversed by particles depends strongly on the zenith angle of their trajectory. For vertical incidence ($\theta=0^{\circ}$), the atmospheric column depth is $X_{\rm total}\simeq 1036\ \text{g/cm}^{2}$, while for horizontal incidence ($\theta=90^{\circ}$), it increases to $X_{\rm total}\simeq 3.65\times 10^{4}\ \text{g/cm}^{2}$. Consequently, the relative probabilities for hadronic interactions and decays, as well as the associated energy losses of secondary particles, can vary significantly with zenith angle. This in turn makes the atmospheric lepton flux dependent on the zenith angle.
The cascade equations can be solved approximately in the hadron-decay-dominated and hadron-interaction-dominated regimes Lipari:1993hd ; Gondolo:1995fq . This method, called the $Z$-moment method, approximates eq. (3) for production and decay processes by putting $\phi_{k}(E^{\prime},X)/\phi_{k}(E,X)\simeq\phi_{k}(E^{\prime},0)/\phi_{k}(E,0)$ so that
$S(k\to j;E,X)\simeq\frac{\phi_{k}(E,X)}{\lambda_{k}(E)}\int_{E}^{\infty}dE^{\prime}\frac{\phi_{k}(E^{\prime},0)}{\phi_{k}(E,0)}\frac{\lambda_{k}(E)}{\lambda_{k}(E^{\prime})}\frac{dn(k\to j;E^{\prime},E)}{dE}=\frac{\phi_{k}(E,X)}{\lambda_{k}(E)}Z_{kj}(E)\,.$ (8)
The atmospheric lepton flux at sea level is written as combinations of $Z$-moments, interaction lengths and lifetimes.
An alternative approach for evaluating the atmospheric lepton fluxes is to solve the cascade equations directly. The Matrix Cascade Equations (MCEq) code fedynitch2015calculationconventionalpromptlepton ; Fedynitch:2018cbl solves the cascade equations using a matrix approach. While the MCEq results are more precise, given hadronic production inputs, the $Z$-moment method is straightforward to implement. The two approaches differ by $5-20\%$ depending on energy and angle Gaisser:2019xlw . In this paper, we use the MCEq code with some modifications outlined below to evaluate all atmospheric lepton fluxes. Another approach is to use Monte Carlo air shower simulations. Recent work LudwigNeste:2025mcf shows good agreement with MCEq.
2.2 Muon and muon neutrino fluxes from MCEq
Figure 2: Results for $E^{3}\phi(E)$ for $\mu^{+}+\mu^{-}$ and $\nu_{\mu}+\bar{\nu}_{\mu}$ for zenith angle $\theta=0^{\circ}$ (vertical), showing the individual contributions of light, heavy and unflavored mesons, as determined by MCEq using the Sibyll-2.3c hadronic interaction model and the H3a cosmic ray flux model.
MCEq is equipped with several hadronic interaction models and cosmic ray nucleon flux models. We use the Sibyll2.3c Riehn:2017mfm ; Fedynitch:2018cbl model for hadronic interactions, and the H3a cosmic ray flux and composition model Gaisser:2011klf . We discuss the impact of different hadronic models and all-nucleon cosmic ray flux inputs in Section 4.3.
The implementation of charm production in Sibyll-2.3c follows a phenomenological model based on soft and hard minijets Fedynitch:2018cbl . For LHCb kinematic distributions for charmed-meson production, the overall combination of perturbative and non-perturbative contributions in Sibyll-2.3c is consistent with the next-to-leading order QCD predictions of refs. Gauld:2015yia ; Garzelli:2016xmx ; PROSA:2015yid that come from collinear factorization with partonic cross sections from perturbative QCD and non-perturbative contributions from both parton distribution functions and fragmentation functions/hadronization. While some components of the implementation of charm production in Sibyll-2.3c are described as “intrinsic” charm, the charm hadrons do not have the same energy distribution, shape and normalization, as in the model for intrinsic charm that we introduce in Section 3. Even though Sibyll-2.3c’s implementation of charm production has a weighted mix of perturbative and non-perturbative components, we will nevertheless refer to Sibyll-2.3c’s contribution to the prompt atmospheric lepton fluxes as coming from perturbative QCD (pQCD($c\bar{c}$)) to distinguish it from the intrinsic charm contribution we introduce in Section 3.
Figure 2 shows the MCEq results for the separate contributions of pions, kaons and (prompt) charm hadrons to the muon (left) and muon neutrino (right) atmospheric fluxes for zenith angle $\theta_{\rm zen}=0^{\circ}$ (vertical). As noted above, the muon flux also includes contributions from light unflavored mesons to its prompt component. The standard release of MCEq does not include $\Lambda_{c}$ production and decay. We include perturbative $\Lambda_{c}$ production by extracting the $D^{0}$ production spectrum from MCEq and applying a constant suppression factor of 0.10 that reflects the relative charm quark fragmentation fractions to $\Lambda_{c}$ compared to $D^{0}$ in proton-proton collisions Lisovyi:2015uqa . The final results of our paper are not very sensitive to this suppression factor because charm contributions to the atmospheric $\mu^{+}+\mu^{-}$, $\nu_{\mu}+\bar{\nu}_{\mu}$ and $\nu_{e}+\bar{\nu}_{e}$ fluxes are dominated by the semi-leptonic decays of $D^{\pm},\ D^{0}$ and $\bar{D}^{0}$ mesons. For semi-leptonic $\Lambda_{c}$ decays, we follow refs. Bugaev:1998bi and Buras:1976dg for the energy distributions of muons and neutrinos, which we have also described in the Appendix. While the details of the semi-leptonic $\Lambda_{c}$ decay distribution are not essential for the pQCD($c\bar{c})$ contributions to the prompt lepton fluxes, a $\Lambda_{c}$ is always produced in our implementation in intrinsic-charm induced events, as discussed in section 3.
Figure 2 exhibits an high-energy behavior that reflects the dominance of the prompt contributions to the atmospheric lepton fluxes, visible in their harder spectra compared to the conventional fluxes. The plots show fluxes for vertical trajectories entering from the zenith, $\theta_{\rm zen}=0^{\circ}$. The zenith angle dependence is primarily in the conventional flux, where pions and kaons are more likely to interact than decay. The conventional flux increases with zenith angle according to $1/\cos\theta_{\rm zen}$ for $\theta_{\rm zen}\leq 70^{\circ}$ Klimushin:2000cy . At very high energies, the prompt contributions will also vary with zenith angle, however, at the energies we consider for the rest of the paper, $E\lesssim 10^{6}$ GeV, the prompt atmospheric lepton fluxes are essentially isotropic.
The strong suppression of the prompt flux relative to the conventional flux in the lower energy range shown in fig. 2 originates from the production and decay properties of the parent particles. For charm baryons, their relatively large mass makes their production less abundant than the production of lighter mesons such as pions and kaons, introducing a significant production suppression factor. In the case of unflavored mesons, their branching fractions to di-muon pairs is small, so even with a relatively large rate of unflavored meson production, their contributions to the muon flux is comparatively small.
We can also obtain di-muon pairs through the process $\gamma\to\mu^{+}\mu^{-}$. There are two possible sources for this channel. The first is from diffuse cosmic rays interacting in the atmosphere and producing high energy photons, which can subsequently convert into muon pairs. However, this contribution is expected to be very small even at the highest $\gamma$-ray energies Illana:2010gh . The second source is the diffuse $\gamma$-ray flux from the Galactic plane. According to ref. LHAASO:2023gne , at around 100 TeV the diffuse Galactic $\gamma$-ray flux is approximately $10^{-5}$ of the diffuse cosmic ray flux. Thus, diffuse Galactic $\gamma$-ray flux will also provide negligible contribution to the overall atmospheric muon flux.
Figure 3: Left: MCEq predictions for angle-averaged (up to zenith angle of $60^{\circ}$) conventional and prompt ($\mu^{+}+\mu^{-}$) flux, compared to the experimentally measured spectrum of high-energy muons using data for $\theta_{\rm zen}\leq 60^{\circ}$ from refs. Soldin:2023lbr ; Soldin:2018vak ; IceCube:2015wro . The unflavored and pQCD($c\bar{c}$) curves are nearly equal in the muon energy range shown in the figure. Right: Comparison of MCEq pQCD($c\bar{c}$) prompt ($\nu_{\mu}+\bar{\nu}_{\mu}$) flux with the upper limit on the prompt flux from the analysis of up-going muon tracks performed in Ref. IceCube:2016umi , which assumed a prompt neutrino energy spectrum with the same shape as in Enberg et al. PhysRevD.78.043005 . The angle-averaged (up to zenith angle of $60^{\circ}$) conventional neutrino flux from MCEq is also shown. The prompt neutrino flux is isotropic in the energy range considered here. The cosmic ray all-nucleon flux model H3a and the hadronic interaction model Sibyll-2.3c are used as input of MCEq.
2.3 Comparisons with IceCube data and upper limits
Figure 3 shows the angle-averaged atmospheric conventional and prompt $\mu^{+}+\mu^{-}$ flux (left panel) and $\nu_{\mu}+\bar{\nu}_{\mu}$ flux (right panel). For the conventional component, we average over zenith angles $\theta_{\rm zen}=0^{\circ}-60^{\circ}$ to avoid nearly horizontal directions where experimental mis-reconstruction of muon events may be significant IceCube:2015wro . In the left panel, the conventional flux, the pQCD(cc¯)($\bar{c}$) contribution, and the unflavored meson contribution are shown separately, along with the total flux (orange curve). The Sibyll-2.3c hadronic interaction model predicts roughly comparable contributions to the prompt atmospheric $\mu^{+}+\mu^{-}$ flux from heavy flavor and light unflavored mesons which makes them difficult to distinguish in the figure. The IceCube measurement of the angle-averaged muon flux for $\theta_{\rm zen}<60^{\circ}$ is shown by the data points with bin-by-bin errorbars IceCube:2015wro . Above $\sim 10^{5}$ GeV, the MCEq prediction underestimates the observed flux by a factor of $\sim 1.5-2$ (see the red points in fig. 6), although the statistical uncertainties are larger as the energy increases. The right panel of fig. 3 shows IceCube’s upper bound on the prompt neutrino flux assuming isotropy PhysRevD.78.043005 , a reliable approximation in the energy range shown in the figure. The IceCube upper limit on the prompt $\nu_{\mu}+\bar{\nu}_{\mu}$ flux still allows room for additional prompt contributions. In what follows, we explore possible additional sources of prompt leptons, i.e. both heavy flavored hadrons generated by mechanisms involving intrinsic charm and light unflavored mesons, and study their interplay in the production of prompt atmospheric muons and muon neutrinos.
3 Intrinsic charm
Constraints on intrinsic charm contributions to the atmospheric neutrino flux have been discussed in the literature (see, e.g., refs. Ostapchenko:2022thy ; Garzelli:2023jlq ; Laha:2016dri ; Halzen:2016thi ). The potential discrepancy between the atmospheric muon measurements and the predictions from MCEq suggest missing contributions to the atmospheric muon flux. One approach is to add intrinsic charm contributions to the prompt atmospheric muon flux, adjusting the normalization of the intrinsic charm contributions to account for the deficit in the theory predictions for muons, then checking for consistency with the experimental upper bound on the flux of atmospheric muon neutrinos.
We consider a phenomenological implementation Garzelli:2023jlq ; Ostapchenko:2022thy of intrinsic charm in which valence-like momentum fractions are transferred to $\bar{D}^{0}$ and $\Lambda_{c}$ in proton-air ($pA$) scattering and to $D^{-}$ and $\Lambda_{c}$ in neutron-air ($nA$) scattering:
$p+A$ $\to\bar{D}^{0}+\Lambda_{c}+X$ (9)
$n+A$ $\to D^{-}+\Lambda_{c}+X\,.$ (10)
From the cascade equation, eq. (5), one can see that the steeply falling cosmic ray flux allows a small fraction of the total charm production cross section from intrinsic charm to have a large contribution to the lepton flux as long as the energy carried by the charm particles is large, as is the case for intrinsic charm.
Following ref. Ostapchenko:2022thy ; Garzelli:2023jlq , we write the differential cross section for proton-air production of charm hadron $h_{c}$ via non-perturbative intrinsic charm as
$\frac{d\sigma_{p-{\rm air}}^{h_{c}{\rm(intr)}}(E_{p},x_{h_{c}})}{dx_{h_{c}}}=w_{\rm intr}^{c}\sigma_{p-{\rm air}}(E_{p})f_{h_{c}}^{\rm(intr)}(x_{h_{c}})\,$, (11)
where $x_{h_{c}}=E_{h_{c}}/E_{p}$, and similarly for neutron constituents of cosmic rays. The parameter $w_{\rm intr}^{c}$, whose value can be extracted by fits to experimental data, determines the overall level of the intrinsic charm contribution relative to the inelastic proton-air cross section. The quantity $f_{h_{c}}^{\rm(intr)}$ is the fragmentation function for the cosmic ray nucleon to produce a charmed baryon or charmed meson $h_{c}$. We assume that $w_{\rm intr}^{c}$ is independent of energy and identical for protons and neutrons.
Figure 4: $f_{h_{c}^{(intr}}$ of $D$ mesons and $\Lambda_{c}$ as a function of $x_{h_{c}}$ (see eq. 11) using the Regge ansatz Kaidalov:1985jg ; Kaidalov:2003wp with either $a_{N}=-0.5$ (Regge1) or $a_{N}=0$ (Regge2), and using the Hobbs, Londergan and Melnitchouk (HLM) approach Hobbs:2013bia (see text for details).
Several theoretical approaches exist to model the intrinsic charm fragmentation, e.g., Brodsky et al. (BHPS) Brodsky:1980pb ; Brodsky:1981se , Regge ansatz Kaidalov:1985jg ; Kaidalov:2003wp and meson-baryon models, e.g., as in ref. Hobbs:2013bia . We use the Regge ansatz as our starting point, and also show results for the meson-baryon model approach of Hobbs, Londergan and Melnitchouk (HLM) Hobbs:2013bia . The Regge fragmentation function (normalized to unity) is
$f_{h_{c}}^{\rm intr}(x)$ $=$ $N_{\rm intr}\,x^{-a_{\psi}}(1-x)^{-a_{\psi}+2(1-a_{N})}$ (12)
$N_{\rm intr}$ $=$ $\frac{\Gamma(-2a_{\psi}+4-2a_{N})}{\Gamma(-a_{\psi}+1)\Gamma(-a_{\psi}+3-2a_{N})}\,.$
The choices for constants used in eq. (12) are $a_{\psi}=-2.0$, $a_{N}=-0.5$ (Regge1) or $a_{N}=0$ (Regge2). We use $a_{N}=-0.5$ (Regge1) as the default for the results shown here, unless otherwise specified. Furthermore, we assume for the Regge ansatz that
$f_{D^{-}}^{\rm(intr)}(x_{D})$ $=$ $f_{\bar{D}^{0}}^{\rm(intr)}(x_{D})$ (13)
$f_{\Lambda_{c}}^{\rm(intr)}(x_{\Lambda})$ $=$ $f_{\bar{D}^{0}}^{\rm(intr)}(1-x_{\Lambda})\,.$ (14)
These functions are shown in fig. 4.
Hobbs, Londergan and Melnitchouk Hobbs:2013bia determined splitting functions in a meson-baryon model, for which the $\bar{D}^{*0}\Lambda_{c}$ dominates for $pA$ interactions. In fig. 4 we show the $\Lambda_{c}$ distribution using $x_{\Lambda_{c}}=1-x_{\bar{D}^{*0}}$. We account for a diminished $\bar{D}^{0}$ momentum fraction from the decay $\bar{D}^{*0}\to D^{0}\pi^{0}$ and $\bar{D}^{*0}\to D^{0}\gamma$ following refs. Cacciari:2003zu ; Cacciari:2005uk . For $nA$ interactions, $D^{*-}\Lambda_{c}$ fluctuations dominate. The decays of $\bar{D}^{*-}\to D^{-}\pi^{0}$, $\bar{D}^{*-}\to D^{-}\gamma$ and $\bar{D}^{*-}\to\bar{D}^{0}\pi^{-}$ have very similar kinematics and yield $D$ meson $x$-distributions that follows the HLM solid line in fig. 4. For both the Regge and HLM approaches, the charmed baryon carries a larger energy fraction than the charmed meson, although the energy balance is much closer in the HLM approach than with the Regge ansatz.
4 Results
4.1 Angle-averaged atmospheric lepton fluxes
As we saw in fig. 3, the MCEq prediction for the atmospheric muon flux lies below the IceCube data. In this section, we begin by introducing intrinsic charm (IC) contributions to both muon and muon neutrino fluxes. The left panel of fig. 5 shows the best-fit intrinsic charm contribution to the angle-averaged ($\theta_{\rm zen}\leq 60^{\circ}$) atmospheric muon flux, as well as the conventional, prompt light unflavored, prompt pQCD($c\bar{c})$ and the total muon flux contributions calculated using MCEq, along with the IceCube muon data IceCube:2015wro . The best-fit value of the $w_{c}^{\rm intr}$ parameter amounts to $w_{c}^{\rm intr}=3.87\times 10^{-3}$ for the default Regge1 intrinsic charm fragmentation model. This value for $w_{c}^{\rm intr}$ implies that $\sim 0.4\%$ of the nucleon-air cross section is due to intrinsic charm. Despite the small value of $w_{c}^{\rm intr}$, the energy distribution of the charm hadrons following eq. 11 is large because the cosmic ray flux falls so rapidly with increasing cosmic ray energy. Figure 6 shows the ratio of the IceCube data (including experimental error bars) to the MCEq prediction with the addition of intrinsic charm with $w_{c}^{\rm intr}=3.87\times 10^{-3}$.
The $1\sigma$ interval for $w_{c}^{\rm intr}$ based on a $\chi^{2}$ fit is $(2.89-4.85)\times 10^{-3}$. Table 1 shows the best fit values and $1\sigma$ intervals for the Regge1, Regge2 and HLM models. All considered models lead to similar ranges of a few $\times 10^{-3}$ for $w_{c}^{\rm intr}$.
Figure 5: As in fig. 3, the angle-averaged atmospheric fluxes with $\theta_{\rm zen}\leq 60^{\circ}$. Left: Including intrinsic charm (IC) contribution with the weight $w_{c}^{\rm intr}=3.87\times 10^{-3}$, the best-fit for the IceCube data IceCube:2015wro . The gray curve shows the sum of conventional and prompt $\mu^{+}+\mu^{-}$ fluxes including intrinsic charm. Right: The prompt atmospheric muon neutrino flux including intrinsic charm contribution with $w^{c}_{\rm intr}=3.87\times 10^{-3}$ and the upper limit on the prompt flux from IceCube IceCube:2016umi . The sum of perturbative and intrinsic charm contributions to the prompt $\nu_{\mu}+\bar{\nu}_{\mu}$ flux is shown with the gray curve.
On the other hand, the right panel of fig. 5 shows that the prompt perturbative plus intrinsic charm contributions to the prompt atmospheric $\nu_{\mu}+\bar{\nu}_{\mu}$ flux significantly overshoot the IceCube upper bound IceCube:2016umi . To satisfy the prompt atmospheric $\nu_{\mu}+\bar{\nu}_{\mu}$ flux constraint from ref. IceCube:2016umi , $w_{c}^{\rm intr}$ must be reduced, which also implies that intrinsic charm alone cannot account for the apparent deficit in the flux of atmospheric muons at high energies. The IceCube upper bound IceCube:2015wro relies on the shape of the prompt flux from Enberg et al. Enberg:2008te as a function of neutrino energy. We observe that this flux has a slightly different shape than the MCEq prediction. We normalize $w_{c}^{\rm intr}$ to the value of the IceCube upper bound at $E_{\nu}=10^{4}$ GeV, where the upper bound has a minimum. Figure 7 shows the intrinsic charm contributions to the atmospheric muon neutrino (left) and muon (right) fluxes for $w_{c}^{\rm intr}=4.65\times 10^{-4}$. Regge2 and HLM fragmentation models have similar best-fit values to Regge1. With $w_{c}^{\rm intr}$ nearly a factor of 10 smaller than the $\mu^{+}+\mu^{-}$ best-fit value, the atmospheric muon flux data remain above the sum of the MCEq flux and this now reduced intrinsic charm contribution, as shown in the right panel of fig. 7. Figure 8 shows the ratio of IceCube muon data with error bars to the muon flux with and without intrinsic charm, as in fig. 6, but here with $w_{c}^{\rm intr}=4.65\times 10^{-4}$. Such a small value of $w_{c}^{\rm intr}$ has little impact on the muon flux ratio.
Table 1: Values of $w_{c}^{\rm intr}$ for intrinsic-charm model-dependent fit to the angle-averaged muon energy spectrum ($\theta_{\rm zen}\leq$ $60^{\circ}$). Best fit values of $w_{c}^{\rm intr}$ along with the $1\,\sigma$ interval are shown for the two Regge ansatz models Kaidalov:1985jg ; Kaidalov:2003wp and for the Hobbs, Londergan and Melnitchouk (HLM) model Hobbs:2013bia .
Model Best fit $\mathbf{{\it w_{c}}^{\rm intr}}(\times 10^{-3})$ $\mathbf{1\,\sigma}$ interval $(\times 10^{-3})$
[…]
HLM 2.99 2.233.742.23-3.74
Figure 6: Ratio of IceCube muon data with experimental error bars IceCube:2015wro to the muon flux from MCEq both with (blue) and without (red) including intrinsic charm contribution with $w_{c}^{\rm intr}=3.87\times 10^{-3}$.
Figure 7: As in fig. 3, angle-averaged with $\theta_{\rm zen}\leq 60^{\circ}$. Left: Comparison of prompt neutrino spectrum ($\nu_{\mu}+\bar{\nu}_{\mu}$) including intrinsic charm contribution with $w_{c}^{\rm intr}=4.65\times 10^{-4}$ with the upper limit from IceCube IceCube:2016umi . Right: Comparison of angle-averaged muon flux ($\mu^{+}+\mu^{-}$) with IceCube data IceCube:2015wro after including intrinsic charm contribution with $w_{c}^{\rm intr}=4.65\times 10^{-4}$.
Figure 8: Ratio of IceCube muon data with experimental error bars IceCube:2015wro to the muon flux from MCEq both with (blue) and without (red) including intrinsic charm contribution with $w_{c}^{\rm intr}=4.65\times 10^{-4}$.
As a next step, we consider the possibility of a normalization uncertainty in the unflavored light meson contribution to the prompt muon flux. This could arise from a higher multiplicity and/or a harder energy spectrum of light meson production in cosmic ray-air collisions relative to the Sibyll-2.3c model. We define a scale factor $\alpha_{\rm unfl}$ that multiplies the Sibyll-2.3c angle-averaged unflavored light meson contribution to the atmospheric muon flux:
$\phi_{\mu}^{\rm unfl}=\alpha_{\rm unfl}\,\phi_{\mu}^{\rm MCEq,unfl}\,.$ (15)
[…]
We note that the shapes and normalizations of the prompt contributions to the atmospheric muon flux from light unflavored hadrons and from charm hadrons from Sibyll-2.3c in MCEq are nearly identical. The flux shapes are determined by the hadronic interactions and the prompt decays. The nearly equal normalization is accidental: many more light unflavored mesons are produced than charm hadrons, however, the light-meson branching fractions to $\mu^{+}+\mu^{-}$ are very small.
Figure 9: Left: The angle-averaged ($\theta_{\rm zen}\leq 60^{\circ}$) muon flux ($\mu^{+}+\mu^{-}$) including an intrinsic charm contribution with $w_{c}^{\rm intr}=4.65\times 10^{-4}$ and scaling the unflavored flux by the best-fit parameter of $\alpha_{\rm unfl}=4.33$. Right: Without any intrinsic charm contribution, the angle-averaged ($\theta_{\rm zen}\leq 60^{\circ}$) muon flux with scaling the unflavored flux by the best-fit parameter of $\alpha_{\rm unfl}=4.80$.
[…]
Model $\mathbf{{\it w_{c}}^{\rm intr}\times 10^{-4}}$ $\mathbf{\alpha_{\rm unfl}}$ $\mathbf{1\sigma}$ interval
[…]
HLM 3.84 4.29 3.31 - 5.28
4.2 Flux ratios
In addition to the angle-averaged atmospheric muon and muon neutrino fluxes, we also show antiparticle-particle ratios, the muon to muon neutrino flux, and a ratio to illustrate the angular dependence of the atmospheric muon flux.
Figure 10: Left: Muon charge ratio without any intrinsic charm, with intrinsic charm and rescaling the unflavored contribution, with all ratios with fluxes angle averaged for $\theta_{\rm zen}\leq 60^{\circ}$. Right: Anti-neutrino to neutrino flux ratio without and with intrinsic charm. The neutrino and anti-neutrino fluxes do not depend on the unflavored contribution. Ratios are of fluxes angle averaged for $\theta_{\rm zen}\leq 60^{\circ}$ and Regge1 intrinsic charm model.
The left panel of fig. 10 shows the ratios of the angle-averaged fluxes $\phi_{\mu^{+}}/\phi_{\mu^{-}}$ for four choices of $(w_{c}^{\rm intr},\alpha_{\rm unfl})$. With each of these four combinations, with or without intrinsic charm, the flux ratios are nearly identical up to $E_{\mu}\sim 3\times 10^{4}$ GeV, where the flux stems predominantly from charged pion and kaon decays. There are more $\mu^{+}$ than $\mu^{-}$ at fixed energy because of the excess of positive pions with respect to negative ones, e.g., more $\pi^{+}$ than $\pi^{-}$, that is traced to the net positive charges of incident cosmic rays. The ratio changes by $\sim 20\%$ from $\phi_{\mu^{+}}/\phi_{\mu^{-}}]\simeq 1.3$ to $\phi_{\mu^{+}}/\phi_{\mu^{-}}]\simeq 1.1$ as $E_{\mu}$ increases by a factor of $\sim 100$. The choices of $(w_{c}^{\rm intr},\alpha_{\rm unfl})$ yield curves between $E_{\mu}\sim 3\times 10^{4}$ GeV and $5\times 10^{6}$ GeV that differ by only $\sim 10\%$. In any case, the muon charges are not distinguished in neutrino telescopes since there are no magnetic fields in detectors.
The ratios of the angle-averaged fluxes $\phi_{\bar{\nu}_{\mu}}/\phi_{\nu_{\mu}}]$ are shown in the right panel of fig. 10. The value of this ratio at $E_{\nu}=10^{3}$ GeV, amounting to $\sim 0.6$, is less than unity because of the same charge excess of $\pi^{+}$ relative to $\pi^{-}$ that yields $\phi_{\mu^{+}}/\phi_{\mu^{-}}]>1$ at the same energy (see discussion in the previous paragraph). The ratio $\phi_{\bar{\nu}_{\mu}}/\phi_{\nu_{\mu}}]$ increases to $\sim 1$ at the energy where the prompt neutrino flux dominates. The light unflavored mesons do not contribute to the neutrino flux so the neutrino flux does not depend on $\alpha_{\rm unfl}$. The ratio curves differ by $\sim 10\%$ for $E_{\nu}=10^{6}$ GeV. This will have little impact on the ($\nu_{\mu}+\bar{\nu}_{\mu}$) interaction rates, especially since the neutrino and antineutrino cross sections differ by less than 5% at this energy Weigel:2024gzh .
Figure 11 shows the ratio of $\phi_{\mu^{+}+\mu^{-}}/\phi_{\nu_{\mu}+\bar{\nu}_{\mu}}$ as a function of energy for different choices for $w_{c}^{\rm intr}$ and $\alpha_{\rm unfl}$. The two lower curves use as input $\alpha_{\rm unfl}=1$ and either $w_{c}^{\rm intr}=0$ or $w_{c}^{\rm intr}=\penalty 10000\ 4.65\times 10^{-4}$. The black curve for $w_{c}^{\rm intr}=0$ tends to a ratio of 2 in the figure since the unflavored and perturbative charm contributions to the muon flux are approximately equal, and the perturbative charm contributions to the neutrino and muon fluxes are nearly equal. For $E\gtrsim 10^{6}$ GeV, including intrinsic charm decreases the ratio by $\sim 18\%$, since intrinsic charm is relatively more important for the neutrino flux than for the muon flux. At the energies where ratios with $\alpha_{\rm unfl}=1$ shown in fig. 11 differ, the astrophysical neutrino flux already dominates the atmospheric neutrino flux. Extracting the ratio in this energy regime will be difficult. The upper two curves in fig. 11 include the best-fit values of $\alpha_{\rm unfl}$ for $w_{c}^{\rm intr}=0$ and $4.65\times 10^{-4}$. The largest values of the ratio $\phi_{\mu^{+}+\mu^{-}}/\phi_{\nu_{\mu}+\bar{\nu}_{\mu}}$ are obtained when the unflavored meson contribution to the prompt $\mu^{+}+\mu^{-}$ flux is increased with no increase in the prompt $\nu_{\mu}+\bar{\nu}_{\mu}$ flux, namely when $w_{c}^{\rm intr}=0$ and $\alpha_{\rm unfl}=4.80$.
Figure 11: Ratio of muon flux to muon neutrino flux with and without intrinsic charm angle averaged for $\theta_{\rm zen}\leq 60^{\circ}$. The red (black) curve has $\alpha_{\rm unfl}=1$, and $w_{c}^{\rm intr}=4.65\times 10^{-4}$ ($w_{c}^{\rm intr}=0$). The green (blue) curve has $w_{c}^{\rm intr}=4.65\times 10^{-4},\ \alpha_{\rm unfl}=4.33$ ($w_{c}^{\rm intr}=0,\ \alpha_{\rm unfl}=4.8$).
Figure 12: Ratio of muon flux per solid angle $\phi_{\mu^{+}+\mu^{-}}(\theta)$ at zenith angles of $0^{\circ}$ and $60^{\circ}$ to the angle-averaged $\theta_{\rm zen}\leq 60^{\circ}$ muon flux per solid angle $\phi_{\mu^{+}+\mu^{-}}$ without any intrinsic charm (labeled MCEq flux), with $w_{c}^{intr}=4.65\times 10^{-4}$ and $\alpha_{\rm unfl}=1$ (with intrinsic charm), and with $w_{c}^{intr}=4.65\times 10^{-4}$ and $\alpha_{\rm unfl}=4.33$ (with scaling of unflavored meson contributions).
Finally, in fig. 12 we show the ratio of the muon flux per solid angle at zenith angles $\theta_{\rm zen}=0^{\circ}$ and $60^{\circ}$ relative to the angle-averaged muon flux. For the flux dominated by pion decays, the flux scales as $\sim 1/\cos\theta$ for the zenith angle range considered here Klimushin:2000cy . With this functional form, the flux ratios are $\phi_{\mu^{+}+\mu^{-}}(0^{\circ})/\phi_{\mu^{+}+\mu^{-}}(\leq 60^{\circ})=0.72$ and $\phi_{\mu^{+}+\mu^{-}}(60^{\circ})/\phi_{\mu^{+}+\mu^{-}}(\leq 60^{\circ})=1.44$, consistent with the curves in fig. 12 for $E_{\mu}\sim 10^{4}$ GeV. The dominance of prompt contributions, whether from perturbative charm, intrinsic charm or light unflavored mesons, leads to an isotropic distribution and a unit ratio at high energy, independent of angle. Figure 12 shows the trend towards prompt dominance as the ratio curves move closer to one. The transition to prompt-dominated atmospheric muon flux as a function of energy is more rapid with $\alpha_{\rm unfl}=4.33$ than with just intrinsic charm or with the default MCEq.
4.3 Uncertainties
There are a number of uncertainties in the MCEq evaluation of the atmospheric leptons fluxes. We discuss them assuming $w_{c}^{\rm intr}=0$. In our MCEq computations, we do not include any seasonal variations in the atmospheric muon and muon neutrino fluxes. The IceCube collaboration has measured a 3.94.63.9-4.6 % seasonal modulation in the lepton flux for energies of 125 GeV - 10 TeV IceCube:2025wtk , a small variation compared to other uncertainties and mostly relevant to the conventional atmospheric lepton contributions.
A potentially larger uncertainty comes from the hadronic interaction modeling. Within the MCEq framework, a range of Monte Carlo event generators for hadronic interactions are available: besides Sibyll-2.3c, EPOS-LHC Pierog:2013ria , QGSJETII04 Ostapchenko:2010vb and DPMJETIII306 Roesler:2000he can also be used for predictions. The three panels of fig.13 show ratios of fluxes using these three models to our default result using Sibyll-2.3c for the angle-average ($\theta_{\rm zen}\leq 60^{\circ}$) for $\mu^{+}+\mu^{-}$. The upper left panel shows the ratios of the prompt fluxes from light unflavored mesons. DPMJETIII306 has a larger proportion of prompt muons from light unflavored mesons than Sibyll-2.3c does for $E_{\mu}\gtrsim 4\times 10^{5}$ GeV, and almost a factor of two larger for $E_{\mu}=10^{6}$ GeV. The EPOS-LHC angle-averaged muon flux from light unflavored mesons nearly equals that from Sibyll-2.3c. The QGSJETII04 results for muons from light unflavored mesons are $\sim 20\%$ of those from Sibyll-2.3c.
With Sibyll-2.3c, we showed that the prompt atmospheric flux of muons from light unflavored mesons is nearly equal to the prompt atmospheric muon flux from charm. The upper right panel of fig. 13 shows the $\mu^{+}+\mu^{-}$ flux ratios for all prompt atmospheric muons (unflavored plus charm). Neither QGSJETII04 nor EPOS-LHC include charm production, so the ratios are consequently reduced. The prompt atmospheric muon flux from DPMJETIII306 does include charm contributions. Including both unflavored mesons and charm, the ratio is $\sim 1.8$ for $E_{\mu}=10^{6}$ GeV and $\sim 0.85$ for $E_{\mu}=10^{4}$ GeV. The lower panel of fig. 13 shows the ratios of the total $\mu^{+}+\mu^{-}$ flux. In the lower energy range of the figure, the ratios are within $\sim 10\%$, while at high energies where the prompt flux dominates (or has no charm contribution as in the cases of QGSJETII04 and EPOS-LHC), the curves deviate. Comparing the muon fluxes of DPMJETIII306 and Sibyll-2.3c, the total flux ratio is $\sim 0.9$ for $E_{\mu}=10^{5}$ GeV, and $\sim 1.3$ for $E_{\mu}=10^{6}$ GeV. A best-fit rescaling of the DPMJETIII306 unflavored meson contribution gives $\alpha_{\rm unfl}=4.84$ (without the inclusion of an intrinsic charm component). Even though the DPMJETIII306 muon flux for $E_{\mu}=10^{6}$ GeV is larger, the fit is dominated by the muon data with smaller errors, $E_{\mu}\lesssim 3\times 10^{5}$ GeV, where the DPMJETIII306 flux is slightly smaller than the Sibyll-2.3c flux, so the value of $\alpha_{\rm unfl}$ is slightly larger for the DPMJETIII306 muon flux than for the Sibyll-2.3c flux.
Figure 13: Ratios of the angle-averaged ($\theta_{\rm zen}\leq 60^{\circ}$) muon fluxes from different interaction models to the muon flux from Sibyll-2.3c, with $w_{c}^{\rm intr}=0$. Left: Unflavored $\mu^{+}+\mu^{-}$ flux. Right: Prompt $\mu^{+}+\mu^{-}$ flux (unflavored plus charm). Bottom: Total $\mu^{+}+\mu^{-}$ flux (conventional, unflavored and charm). Note that QGSJETII04 and EPOS-LHC do not include charm production.
Figure 14 shows the ratios of the angle-averaged atmospheric total flux (left) and prompt flux (right) for $\nu_{\mu}+\bar{\nu}_{\mu}$ for different interaction models. There are no unflavored contributions to the prompt $\nu_{\mu}+\bar{\nu}_{\mu}$ flux. For $E_{\nu_{\mu}}=10^{4}$ GeV, the interaction model predictions agree to within approximately 10%, while again, DPMJETIII306 has relatively more prompt contributions to the $\nu_{\mu}+\bar{\nu}_{\mu}$ flux than Sibyll-2.3c.
Figure 14: The ratios of the angle-averaged $\nu_{\mu}+\bar{\nu}_{\mu}$ fluxes ($\theta_{\rm zen}\leq 60^{\circ}$) from different interaction models to the $\nu_{\mu}+\bar{\nu}_{\mu}$ flux from Sibyll-2.3c, all with $w_{c}^{\rm intr}=0$. Left: Total $\nu_{\mu}+\overline{\nu}_{\mu}$ flux. Right: Prompt $\nu_{\mu}+\overline{\nu}_{\mu}$ flux. Note that QGSJETII04 and EPOS-LHC do not include charm production.
In fig. 15, we show the ratio of the fluxes evaluated with the H4a Gaisser:2013bla (sometimes called H3p, for instance in ref. IceCube:2016umi ) cosmic ray model relative to the H3a model for muons (left) and muon neutrinos (right). Both H3a and H4a are implemented in MCEq as 3-Peters cycle Peters:1961mxb compositions with five mass groups. However, H4a assumes a composition that becomes proton dominated above the ankle, whereas H3a retains a mixed composition. The H4a model predicts a larger high-energy lepton flux, leading to an enhancement of approximately 35% for muons and 40% for muon neutrinos at $E=5\times 10^{6}$ GeV. However, this results in only modest changes to the fitted intrinsic charm and unflavored scaling parameters. For the Regge1 intrinsic charm model, the best fit values shift to $w_{c}^{\rm intr}=4.35\times 10^{-4},\ \alpha_{\rm unfl}=1$ and $w_{c}^{\rm intr}=0,\ \alpha_{\rm unfl}=3.85$ when using H4a instead of H3a, compared to the values reported in table 2.
Figure 15: The ratio of atmospheric lepton fluxes using as input the H4a all-nucleon primary cosmic ray flux model and the H3a all-nucleon cosmic ray flux model. Left: Angle-averaged ($\theta_{\rm zen}\leq 60^{\circ}$) total $\mu^{+}+\mu^{-}$ flux ratio. Right: Angle-averaged ($\theta_{\rm zen}\leq 60^{\circ}$) total $\nu_{\mu}+\bar{\nu}_{\mu}$ flux ratio.
In ref. Yanez:2023lsy , Yañez and Fedynitch tune a data-driven model (daemonflux) that includes hadronic interaction and cosmic ray flux uncertainties. The tune is calibrated to atmospheric muon flux data, primarily below $E_{\mu}=10^{4}$ GeV. They compare the resulting neutrino flux from daemonflux with the atmospheric neutrino fluxes from Sibyll-2.3d, which yields nearly identical inclusive atmospheric lepton fluxes as Sibyll-2.3c Riehn:2019jet . The Sibyll-2.3d muon neutrino flux is within approximately $\pm 5\%$ of the daemonflux muon neutrino flux for $E_{\nu_{\mu}}=10^{4}-10^{6}$ GeV. There is a little more variation in the electron neutrino flux, with the Sibyll-2.3d electron neutrino flux approximately $15\%$ higher for $E_{\nu_{e}}=10^{4}$ GeV, and closer to 5% higher for $E_{\nu_{e}}=10^{5}$ GeV. Our conclusion is that using a data-driven model such as daemonflux would not change our estimates about uncertainties in the atmospheric muon flux predictions, so there remains an apparent deficit in atmospheric muon flux predictions at high energy.
Different intrinsic charm models can, in principle, yield widely varying atmospheric flux predictions. We have assumed a functional form for the differential cross section for charm hadron production via the intrinsic charm mechanism in nucleon-air interactions according to eq. 11. The resulting contributions from intrinsic-charm-produced $D^{-},\ \bar{D}^{0}$, and $\Lambda_{c}$ to the muon and muon neutrino flux have essentially the same shape. The primary effect of changing the fragmentation model is a shift in the overall normalization of the intrinsic charm contribution, which is reflected in the fitted value of $w_{c}^{\rm intr}$, as shown in tables 1 and 2. The reason the fitted $w_{c}^{\rm intr}$ values differ among the models is that the Regge models produces a larger fraction of $\Lambda_{c}$ baryons at higher energies compared to $D$ mesons, whereas the HLM model yields almost equal distributions of energetic $D$ and $\Lambda_{c}$ hadrons. Since $D$ mesons have higher branching ratios to muon and muon neutrino than $\Lambda_{c}$, and the leptons from $D$ meson decays can emerge with higher energy fractions than for $\Lambda_{c}$ decays (see figs. 18 and 19 in the Appendix), the HLM model gives a relatively larger intrinsic charm contribution to the muon and muon neutrino fluxes, so the best fit normalization $w_{c}^{\rm intr}$ is smaller for HLM fragmentation compared to the Regge models. In other words, the physical differences between the models manifest primarily as a change in the fitted intrinsic charm normalization, rather than in the shape of the resulting lepton spectra. Overall, whether $w_{c}^{\rm intr}$ is fitted to the muon flux or the muon neutrino flux, all three valence-like fragmentation models for intrinsic charm give values for $w_{c}^{\rm intr}$ that are roughly comparable among each other.
5 Discussion and Conclusions
IceCube’s measurements of the $\mu^{+}+\mu^{-}$ atmospheric flux at high energies suggest a missing prompt component. We have focused on fluxes at zenith angles $\theta\leq 60^{\circ}$ to avoid issues with potential mis-reconstruction of muon events in the experimental data IceCube:2015wro . Since the prompt $\nu_{\mu}+\bar{\nu}_{\mu}$ and $\mu^{+}+\mu^{-}$ fluxes from charm are nearly equal, an increase of the prompt $\mu^{+}+\mu^{-}$ flux from charm to improve agreement with the high-energy muon data (which would correspond to taking $w_{c}^{\rm intr}=3.87\times 10^{-3}$) can violate the upper bounds on the prompt $\nu_{\mu}+\bar{\nu}_{\mu}$ flux. Indeed, in addition, such a high value of $w_{c}^{\rm intr}$ would lead to an excess of very forward production of $\Lambda_{c}^{+}$ in low-energy collisions since we have assumed that $w_{c}^{\rm intr}$ is energy independent. The value of $w_{c}^{\rm intr}=4.65\times 10^{-4}$ is sufficiently small that the intrinsic charm cross section is less than 5% of the perturbative cross section for fixed-target $pp$ production of $c\bar{c}$ at $E_{p}=10^{5}$ GeV and less than $\sim 1\%$ for $E_{p}\gtrsim 10^{6}$ GeV, well within the next-to-leading order perturbative Nason:1987xz ; Bhattacharya:2016jce ; Garzelli:2015psa and measurement PHENIX:2006tli ; HERA-B:2007rfd ; ALICE:2012inj ; STAR:2012nbd ; ATLAS:2015igt ; LHCb:2013xam uncertainties for open charm production. However, this smaller value of $w_{c}^{\rm intr}$ still underpredicts the atmospheric $\mu^{+}+\mu^{-}$ flux at high energy, as shown in fig. 7.
Figure 16: Left: Angle-averaged (up-to $\theta_{\rm zen}\leq 60^{\circ}$) electron neutrino flux ($\nu_{e}+\bar{\nu}_{e}$) scaled by $E_{\nu_{e}}^{3}$ with the inclusion of intrinsic charm with $w_{c}^{\rm intr}$ values of $3.87\times 10^{-3}$ (dashed lines) and $4.65\times 10^{-4}$ (solid lines). Right: Angle-averaged for full zenith angle range, the $\nu_{e}+\bar{\nu}_{e}$ flux contributions scaled by $E_{\nu_{e}}^{2}$, along with a data point from IceCube IceCube:2015mgt .
We have used IceCube’s upper bound on the prompt atmospheric $\nu_{\mu}+\bar{\nu}_{\mu}$ flux from measurements of through-going muons IceCube:2016umi to constrain the intrinsic charm contribution. IceCube’s upper bound relies on the Enberg et al. (ERS) Enberg:2008te prompt neutrino flux, modified by an input H3p cosmic ray flux and a single power law model of the diffuse astrophysical neutrino flux. A recent re-evaluation of the IceCube sensitivity from muon tracks, using H4a and GST all-nucleon cosmic ray fluxes and a single power law for the astrophysical flux, yields an 90% CL upper bound on the prompt flux normalization of 2.9 times the Sibyll-2.3c prompt neutrino flux IceCube:2023mrq . With either cascades or northern tracks and a single power law astrophysical flux, the 90% CL upper bound on the prompt flux normalization is 1.34 and 3.02 times the Sibyll-2.3c prompt neutrino flux, respectively Bottcher:2025hcz . Upper limits vary with the assumed astrophysical neutrino flux, and whether or not combined fits are made. Despite the model dependence of the upper bound on the prompt atmospheric $\nu_{\mu}+\bar{\nu}_{\mu}$ flux, we conclude that there is room for only a small intrinsic charm contribution to the prompt neutrino flux.
Measurements of the high-energy atmospheric $\nu_{e}+\bar{\nu}_{e}$ flux may shed light on the prompt flux contribution. While the prompt atmospheric flux of $\nu_{e}+\bar{\nu}_{e}$ is very nearly equal to the prompt atmospheric $\nu_{\mu}+\bar{\nu}_{\mu}$ flux, the conventional fluxes have larger differences. The conventional atmospheric $\nu_{e}+\bar{\nu}_{e}$ flux is about a factor of 20 lower than the conventional $\nu_{\mu}+\bar{\nu}_{\mu}$ flux because nearly 100% decays of $\pi^{+}$ lead to $\nu_{\mu}+\mu^{+}$, and at the energies under consideration here, the muons do not decay. Kaon production and decay dominate as the source of conventional atmospheric $\nu_{e}+\bar{\nu}_{e}$ flux. Consequently, the prompt $\nu_{e}+\bar{\nu}_{e}$ flux becomes larger than the corresponding conventional flux at energies lower than in the case of $\nu_{\mu}+\bar{\nu}_{\mu}$. This is illustrated in fig. 16, where we show both pQCD and two versions of the prompt flux, both with and without including intrinsic charm. In the left panel of fig. 16, the $\nu_{e}+\bar{\nu}_{e}$ flux is scaled by $E_{\nu_{e}}^{3}$ (as with the prior flux figures). As already noted, the dashed lines that include $w_{c}^{\rm intr}=3.87\times 10^{-3}$ violate the upper limit for the prompt $\nu_{\mu}+\bar{\nu}_{\mu}$ flux.
The angle-averaged atmospheric electron neutrino flux was measured by DeepCore/IceCube for $E_{\nu}=80$ GeV $\penalty 10000\ -\penalty 10000\ 6$ TeV IceCube:2012jwm and up to a highest energy bin at $\log_{10}(E_{\nu}/{\rm GeV})=4.0-5.0$ by IceCube IceCube:2015mgt . The right panel of fig. 16 shows the fluxes in the left panel, now scaled by $E_{\nu_{e}}^{2}$, and the data point of the highest energy bin of the IceCube atmospheric electron neutrino flux measurement IceCube:2015mgt . The measurements of the atmospheric $\nu_{e}+\bar{\nu}_{e}$ flux is difficult at high energy, in part because the diffuse astrophysical neutrino flux is large. For reference, the IceCube’s high-energy starting event measurements of the per-flavor astrophysical flux in the energy region of $E_{\nu}=100$ TeV, assuming a single power law, is $E_{\nu}^{2}\phi_{\nu+\bar{\nu}}^{\rm per\ flavor}\sim 2\times 10^{-8}(E_{\nu}/100\ {\rm TeV})^{-2.5}$ IceCube:2020acn ; IceCube:2024fxo ; IceCube:2025dlr . To the extent that the diffuse astrophysical neutrino flux is inferred from $\nu_{\mu}+\bar{\nu}_{\mu}$ measurements, it may be possible to understand the relative contributions of the atmospheric and prompt $\nu_{e}+\bar{\nu}_{e}$ fluxes in the measurements.
Figure 17: Angle-averaged ($\theta_{\rm zen}\leq 60^{\circ}$) atmospheric $\phi_{\mu^{+}+\mu^{-}}$ scaled by $E_{\mu}^{3.7}$ with $w_{c}^{\rm intr}=0$ and $\alpha=1,\ 3.81,\ 4.80,\ 5.78$ and $8.60$. The IceCube data from ref. IceCube:2015wro are also shown.
The IceCube limits on the prompt atmospheric $\nu_{\mu}+\bar{\nu}_{\mu}$ flux put significant constraints on the possibility that intrinsic charm can account for IceCube’s high energy $\mu^{+}+\mu^{-}$ measurements, albeit with large errorbars. The measurements suggest a prompt component that could otherwise come from prompt decays of light unflavored mesons, as we discussed in section 4 with an enhancement factor of $\alpha_{\rm unfl}=4.80$ larger than the MCEq default contribution. The IceCube Collaboration reports that the combined heavy-flavor prompt plus unflavored prompt contributions to the atmospheric muon flux correspond to 4.75 times the ERS flux with a 90% CL range of 2.33–9.34 times the ERS flux, assuming the H3a all-nucleon cosmic ray energy spectrum IceCube:2015wro . Their conclusion is consistent with their angular analysis as well. IceCube’s results are consistent with the enhancements of the prompt contributions to the muon flux discussed here.
The large experimental error bars at high energies make conclusive statements about the role of prompt contributions to the atmospheric $\mu^{+}+\mu^{-}$ flux difficult. The spectral shapes of the prompt contributions from charm and from unflavored mesons to the $\mu^{+}+\mu^{-}$ flux are nearly identical in the energy range of interest, so we can trade enhancements of the charm contributions for enhancements of the light unflavored contributions.
We again show the large experimental errors in the measured $\mu^{+}+\mu^{-}$ flux scaled by $E_{\mu}^{3.7}$ in fig. 17, where the experimental data are compared to theory predictions for different values of $\alpha_{\rm unfl}$ and $w^{c}_{\rm intr}=0$. In particular, the figure shows curves for $\alpha_{\rm unfl}=1$, and $\alpha_{\rm unfl}=4.8^{+0.98}_{-0.99}$, corresponding to the $\chi^{2}$ best-fit value with $1\sigma$ errors. We also show a curve for $\alpha=8.6$, to show a more extreme scaling. It is still consistent within $2\sigma$ uncertainties. All of the scaling factors multiply the unflavored meson contributions to the prompt flux according to interactions modeled by Sibyll-2.3c. As the lower panel of fig. 13 shows, DPMJETIII306 would lead to a larger high-energy contribution. However, its energy dependence cannot account for the apparent excess of atmospheric muons in the hundreds of TeV energy range.
The extraction of the prompt atmospheric muon and neutrino fluxes from neutrino telescope data remains a topic of tremendous interest. In combination, SHiP Alekhin:2015byh ; Bai:2018xum and the forward physics program Anchordoqui:2021ghd ; Feng:2022inv with neutrinos Bai:2022xad ; Bai:2022jcs and studies of unflavored meson production at accelerator experiments, in combination with a better modeling of the unflavored component in Monte Carlo event generators for (astro)-particle physics and more precise tunings of the relevant parameters of the latter Albrecht:2025kbb , will be important to better interpret existing and future data, potentially resolving the tension between current theoretical predictions and measurements of the high-energy atmospheric $\mu^{+}+\mu^{-}$ flux.
Acknowledgements.
We are grateful to Anatoli Fedynitch for help and tips on the use of the MCEq framework, to Felix Riehn for discussions on the physics implemented in the Sybill event generator and comparisons with our QCD predictions and to Sergey Ostapchenko and Johannes Blümlein for sharing with us their considerations on intrinsic charm. The model options for intrinsic charm used in this work were suggested by Sergey Ostapchenko. This work was supported in part by the US Department of Energy under grant DE-SC-0010113, by the Deutsche Forschung Gemeinschaft within the FOR2926 Resarch Unit (Project ID 40824754) and under Germany’s Excellence Strategy – EXC 2121 “Quantum Universe” – 390833306. For facilitating portions of this research, DG wishes to acknowledge the Center for Theoretical Underground Physics and Related Areas (CETUP*), The Institute for Underground Science at Sanford Underground Research Facility (SURF), and the South Dakota Science and Technology Authority for hospitality and financial support, as well as for providing a stimulating environment.
Appendix: Three-body decays of charm mesons and baryons
The energy distributions of leptons from charm meson semi-leptonic decays such as $D^{0}\to K^{-}\mu^{+}\nu_{\mu}$ are often evaluated using the matrix element squared for the charm quark decay. For the spin-summed and spin-averaged matrix element squared for the decay $c\to s\mu^{+}\nu_{\mu}$ one obtains
$\overline{|M|^{2}}=64\,G_{F}^{2}|U_{cs}|^{2}(p_{s}\cdot p_{\nu_{\mu}})\,(p_{\bar{\mu}}\cdot p_{c})=64\,G_{F}^{2}|U_{cs}|^{2}m_{c}E_{\mu}p_{\nu}(m_{c}-E_{\mu}+p_{\mu}\cos\theta_{\mu\nu})\,$, (16)
where the last expression is obtained by evaluating the 4-vector scalar products and employing energy-momentum conservation in the rest frame of the charm quark, and $U_{cs}$ is the quark dependent part of the matrix element. Here, $m_{c}$ is the charm quark mass, $E_{\mu}$ and $p_{\mu}$ are the $\mu^{+}$ energy and total momentum, respectively, $p_{\nu}$ is the magnitude of the neutrino 3-momentum and $\theta_{\mu\nu}$ is the angle between the anti-muon and neutrino momenta.
Integrating out the strange quark momentum, integrating over $\cos\theta_{\mu\nu}$ which eliminates the delta function associated to energy conservation, and integrating over the remaining azimuthal and solid angles of ${\bf p}_{\mu}$ and ${\bf p}_{\nu}$ we obtain
$d\Gamma=16\,\pi^{2}G_{F}^{2}|U_{cs}|^{2}\int dE_{\mu}\int dp_{\nu}\,E_{\mu}(m^{2}-m_{c}E_{\mu})\,$, (17)
with $m^{2}\equiv(m_{c}^{2}+m_{\mu}^{2}-m_{s}^{2})/2$. We define $\mu_{s}\equiv m_{s}^{2}/m_{c}^{2}$ and $\mu_{\mu}\equiv m_{\mu}^{2}/m_{c}^{2}$.
We illustrate the impact of final state masses on the $\mu^{+}$ energy distribution in the charm rest frame and in the frame where the charm particle has energy $E_{c}$. In the charm rest frame, since we are interested in the $\mu^{+}$ energy distribution, one has to perform a final integration over the neutrino momentum within its kinematic limits. Further defining $x\equiv 2E_{\mu}/m_{c}$, this yields
$\frac{d\Gamma}{dx}$ $=$ $\pi^{2}G_{F}^{2}m_{c}^{5}\frac{\left[{x^{2}}-4\mu_{\mu}\right]^{1/2}x\left[1+\mu_{\mu}-\mu_{s}-x\right]^{2}}{1+\mu_{\mu}-x}\,$, (18)
where $x$ is constrained to the interval
$x_{\rm min}\equiv\frac{2m_{\mu}}{m_{c}}\leq x\leq 1+\mu_{\mu}-\mu_{s}\equiv x_{\rm max}\,.$ (19)
In the limit $m_{\mu}=0$ this leads to
$\frac{d\Gamma}{dx}=\,\pi^{2}G_{F}^{2}m_{c}^{5}\frac{x^{2}\left[1-\mu_{s}-x\right]^{2}}{1-x}\,$, (20)
for $0\leq x\leq 1-\mu_{s}$, consistent with the literature. We note that in charm decay, the $\mu^{+}$ is the leptonic antiparticle, while in muon decay $\mu\to\nu_{\mu}e\bar{\nu}_{e}$, the corresponding antiparticle is the $\bar{\nu}_{e}$.
Finally, one has to perform a Lorentz boost with a Lorentz factor $\gamma$, corresponding to velocity $\beta$, to the laboratory frame,
$E_{\mu}=\gamma E_{\mu}^{0}\left(1+\beta\frac{p_{\mu}^{0}}{E_{\mu}^{0}}\cos\theta\right)\,$, (21)
where the quantities with superscript 0 now refer to the charm quark rest frame and thus to the $\mu^{+}$ energy and momentum appearing in the previous equations. Defining $y\equiv E_{\mu}/E_{c}$, where $E_{c}=\Gamma m_{c}$ is now the charm quark energy in the laboratory frame, one obtains
$y=\frac{x}{2}+\frac{\beta}{2}\left[{x^{2}}-4\mu_{\mu}\right]^{1/2}\cos\theta\,$, (22)
and
$\frac{d\Gamma}{dy}=\int_{x_{\rm min}(y)}^{x_{\rm max}(y)}\frac{dx}{2}[{x^{2}}-4\mu_{\mu}]^{-1/2}\frac{d\Gamma}{dx}\,$, (23)
where $d\Gamma/dx$ is given by eq. (18). Here, the kinematic range of $y$ is given by
$\frac{x_{\rm max}}{2}-\frac{\beta}{2}\left[{x^{2}}-4\mu_{\mu}\right]^{1/2}\leq y\leq\frac{x_{\rm max}}{2}+\frac{\beta}{2}\left[{x^{2}}-4\mu_{\mu}\right]^{1/2}\,$, (24)
and the integration boundaries in eq. (23) are given by
$x_{\rm min}(y)=\max(x_{\rm min},x_{-})\,,\quad x_{\rm max}(y)=\min(x_{\rm max},x_{+})\,$, (25)
where $x_{\rm min}$ and $x_{\rm max}$ are given by eq. (19) and
$x_{\pm}\equiv 2\gamma^{2}\left[y\pm\left(y^{2}-\frac{y^{2}+\beta^{2}\mu_{\mu}}{\gamma^{2}}\right)^{1/2}\right]\,.$ (26)
As long as the matrix element has the structure of eq. (16), the above equations can be applied to any other 3-body decays by substituting $m_{c}$ and $m_{s}$ with the relevant particle masses.
Figure 18: The center-of-mass frame $x=2E_{\mu}^{0}/m_{D}$ (left panel) and the boosted $y=E_{\mu}/E_{D}$ (right panel) distributions of $\mu^{+}$ and $\nu_{\mu}$ from $D^{-}$ decay, in the case of massive muon.
Figure 18 shows the impact of the muon mass on the $\mu^{+}$ normalized energy distribution in the charm meson rest frame where $x=2E_{\mu}^{0}/m_{D}$ (left) and in the lab frame where the charm meson has energy $E_{c}$ and $y=E_{\mu}/E_{D}$ (right). A similar procedure can be applied to the $\nu_{\mu}$ energy distribution. In Fig. 18, we show the $\mu^{+}$ and $\nu_{\mu}$ distributions in $x$ and $y$ from $D^{0}\to K^{-}\mu^{+}\nu_{\mu}$.
Figure 19: Rest frame $x=2E_{\ell}^{0}/(M_{\Lambda_{c}})$ (left panel) and boosted frame $y=E_{\ell}/E_{\Lambda_{c}}$ (right panel) distributions of $\mu^{+}$ and $\nu_{\mu}$ from $\Lambda_{c}$ decays.
Table 3: The comparison of $Z_{h\mu}^{\rm scaling}$ and $Z_{h\nu_{\mu}}^{\rm scaling}$ for the decay of $h\equiv D^{\pm},\,\Lambda_{c}^{+}$ to muons and muon neutrinos in the scaling approximation where the cosmic ray flux energy spectrum is $\sim E^{-\gamma}$.
$h$ Spectral Index $\gamma$ $Z_{h\nu}^{\rm scaling}/Z_{h\mu}^{\rm scaling}$
$D^{\pm}$ 3 1.15
2.7 1.13
$\Lambda_{c}^{+}$ 3 1.16
2.7 1.14
Traditionally, the $\mu^{+}$ energy distribution from $\Lambda_{c}$ decays is evaluated using three decay form factors Buras:1976dg ; Pietschmann:1984en . (See also the appendix of the preprint version of ref. Bugaev:1998bi .) We evaluated the energy distributions for $\mu^{+}$ and $\nu_{\mu}$ in an effective three-body semi-leptonic decay including the muon mass and an effective final-state strange hadron mass $m_{s}^{\rm eff}=1.27$ GeV Bugaev:1998bi . The $\mu^{+}$ and $\nu_{\mu}$ $x$-distribution in the rest frame and $y$-distribution in the boosted frame are shown in the left and right panel of fig. 19.
It was noted already in Ref. Illana:2010gh ; Volkova:2011zza that, due to the VAV-A couplings of leptons in the weak charged current, neutrinos carry about 15% more of the parent charm hadrons’ momentum than the muon. We confirm this with our distributions in the $Z$-moment approximation. In this approximation, the lepton flux is proportional to $Z_{h\ell}$, for hadron $h$ and lepton $\ell=\mu,\nu$, where
$Z_{h\ell}(E)=\int_{E}^{\infty}dE_{h}\Biggl(\frac{\phi_{h}(E_{h})}{\phi_{h}(E)}\Biggr)\frac{d_{h}(E)}{d_{h}(E_{h})}\frac{1}{\Gamma(h)}\frac{d\Gamma(h\to\ell)}{dE}\,$, (27)
in terms of the lepton energy $E$, hadron energy $E_{h}$, the flux of hadrons $\phi_{h}$ that depends on energy and the distance traveled in a mean $h$ decay time $d_{h}(E)\propto E$. For an incident cosmic ray flux that has power law scaling $\phi_{\rm CR}\propto E^{-\gamma}$, charm hadrons in the prompt regime scale as $\phi_{h}\propto E^{-\gamma+1}$ Lipari:1993hd ; Gondolo:1995fq . The decay $Z$-moment is energy independent in this approximation and can be written as
$Z_{h\ell}^{\rm scaling}=\int_{0}^{1}dy\,y^{\gamma-1}\frac{1}{\Gamma(h)}\frac{d\Gamma(h\to\ell)}{dy}\,.$ (28)
Table 3 shows the ratio $Z_{h\nu}^{\rm scaling}/Z_{h\mu}^{\rm scaling}$ for $D^{\pm}$ and $\Lambda_{c}^{+}$ semileptonic decays, for $\gamma=2.7$ and $\gamma=3$ in a scaling cosmic ray energy spectrum $\sim E^{-\gamma}$. In the case of $D^{0}$ and $\bar{D}^{0}$ decays the ratio remains the same as for $D^{\pm}$, since both use eq. (23).